OpenStudy (anonymous):

The map shows the route for a race. You are at A, which is 6000 feet from the first checkpoint B. The second checkpoint C is located at the midpoint between B and the end of the race D. The total race is 3 miles. How far apart are the 2 checkpoints?

OpenStudy (anonymous):

 AD=15840 AD-AB=BD=15840-6000=9840 BC=9840/2=4920 4920 ft apart